how many grams of O2 will react completely with 35.0 grams PH3?



1
Expert's answer
2021-11-15T10:55:08-0500

2PH3 + 4O2 → P2O5 + 3H2O

Molar mass of PH3= 33.99758 g/mol

35/33.99758 = 1.029mol

Molar mass of O2= 31.998

Mole ratio= 2:4

= 2×1.029×31.998= 65.883g